OpenStudy (cold3693):

Solve each equation: x^4 - x = 0 I got as far as x(x^3 - 1)=0

geerky42 (geerky42):

HINT: \(a^3-b^3 = (a-b)(a^2+ab+b^2)\) You have a = x and b = 1